Determine the total solar energy flux absorbed by the Earth (Eabs) (albedo a = 0.36, and planetary radius r = 6371 km) (in Watts).
Answer #1

Thank you for asking. Let's solve this.

Albedo= 0.36

Radius= 6371km= 6,371,000 meters

We know that solar "insolation" constant is given by - 1,361 watts per square meter

Therefore, total solar energy flux absorbed by the earth = Solar constant * (1- albedo) * R2

=> Energyabsorbed = 1,361*(1-0.36)* 3.14* (6371000)2

= 1,361* 0.64* 3.14* 40,589,641*106

= 871.04*3.14* 40,589,641*106

= 111,015,330,815.44*106 watts

= 111.015*1015 watts

I hope this will help you. In case if you have any questions, please feel free to ask. I'll be more than happy to explain it for you. Thanks again.